A man is 5years older than his wife who's daughter is half her age. How old re they if their ages sum up to a full percentage?
Answer by Reema Kumari(5)

Sol:-Let the age of daughter be x.
then,mother's age=2x
and,father's age=2x+5
A/Q
x+2x+2x+5=100
5x+5=100
5(x+1)=100
x+1=20
x=19
Since,daughter's age=19yrs
mother's age=2x=2*19=38yrs
father's age=2x+5=38+5=43yrs